State suitable conditions and prove that (𝑓𝑔) β€² = 𝑓𝑔 β€² + 𝑓 ′𝑔.


1
Expert's answer
2021-12-16T10:00:48-0500

Let us state the suitable conditions: if the functions "f" and "g" are differentiable in each point of their common domain, then the function "fg" is differentiable in each point of this domain and "(\ud835\udc53\ud835\udc54)'= \ud835\udc53\ud835\udc54' + \ud835\udc53'\ud835\udc54."


Let us prove that "(\ud835\udc53\ud835\udc54)'= \ud835\udc53\ud835\udc54' + \ud835\udc53'\ud835\udc54:"


"(f(x)g(x))'
=\lim\limits_{\Delta x\to 0}\frac{f(x+\Delta x)g(x+\Delta x)-f(x)g(x)}{\Delta x}
\\=\lim\limits_{\Delta x\to 0}\frac{f(x+\Delta x)g(x+\Delta x)-f(x+\Delta x)g(x)+f(x+\Delta x)g(x)-f(x)g(x)}{\Delta x}
\\=\lim\limits_{\Delta x\to 0}\frac{f(x+\Delta x)[g(x+\Delta x)-g(x)]+[f(x+\Delta x)-f(x)]g(x)}{\Delta x}
\\=\lim\limits_{\Delta x\to 0}\frac{f(x+\Delta x)[g(x+\Delta x)-g(x)]}
{\Delta x}+ \lim\limits_{\Delta x\to 0}\frac{[f(x+\Delta x)-f(x)]g(x)}{\Delta x}

\\=\lim\limits_{\Delta x\to 0}f(x+\Delta x)\lim\limits_{\Delta x\to 0}\frac{g(x+\Delta x)-g(x)}
{\Delta x}+ \lim\limits_{\Delta x\to 0}\frac{f(x+\Delta x)-f(x)}{\Delta x}g(x)
\\=f(x)g'(x)+f'(x)g(x)."